OpenStudy (anonymous):

Graph the line that passes through the point (6, -8) and has slope -1/3.

OpenStudy (anonymous):

y+8=-1/3(x-6) y=-1/3x -6

OpenStudy (anonymous):

1st you need to find the equation of that line and fortunately, we are given everything we need to find an equation of that line. gradient is -1/3 and point is (6;-8). therefore: see the solution for finding the equation first: